# A Single-Loop Gradient Algorithm for Pessimistic Bilevel Optimization via Smooth Approximation

### v1

@@ 1 Introduction
where $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$ represent the upper-level and lower-level decision variables, respectively, and ${X}\subseteq\mathbb{R}^{n}$ and ${Y}\subseteq\mathbb{R}^{m}$ are closed convex sets. The functions $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ and $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are the upper-level and lower-level objective functions, respectively. Bilevel optimization naturally models non-cooperative game between two players, often referred to as a Stackelberg game vonStackelbergHeinrich1953TTot . When the lower‐level problem in bilevel optimization admits multiple optimal solutions for a given $x$, the corresponding decision variable $y$ in the upper‐level objective becomes ambiguous. To resolve this, bilevel optimization is commonly formulated in two distinct settings: Optimistic Bilevel Optimization (OBO) and Pessimistic Bilevel Optimization (PBO).
@@
In recent years, bilevel optimization has garnered substantial interest within the machine learning community, finding applications in areas such as hyperparameter optimization franceschi2017bridge , adversarial learning zhang2022revisiting , reinforcement learning zheng2024safe , and meta-learning gao2020modeling , among others, where first-order gradient-based methods are preferred for their efficiency and scalabilitybottou2018optimization ; lin2020accelerated ; sra2011optimization .
Much of the existing bilevel research focuses on the OBO case, for which numerous fully first-order gradient-based algorithms suitable for large-scale machine learning tasks have been developed kwon2023fully ; liu2022bome ; liu2024moreau ; shen2023penalty , often by reformulating the problem using Karush-Kuhn-Tucker (KKT) conditions or through value-function-based constraints. While OBO benefits from a well-established algorithmic toolkit, PBO remains comparatively underexplored from an algorithmic standpoint. PBO offers a robust framework for leaders concerned with worst-case follower responses and a growing body of work has explored the potential of PBO in various machine learning applications such as adversarial learning bruckner2011stackelberg ; benfield2024classification , contextual optimization bucareydecision ; jimenez2025pessimistic and hyperparameter optimization ustun2024hyperparameter . Outside the machine learning domain, PBO has found applications in many other practical scenarios, including but not limited to demand response managementkis2021optimistic , rank pricing and second-best toll pricing calvete2024novel ; ban2009risk , production-distribution planning zheng2016pessimistic , and gene knockout modelzeng2020practical . Yet, the inherent max-structure at the upper level of PBO creates a more complex, three-level-like structure (min-max-min), making the direct application of gradient-based techniques developed for OBO challenging. Although several PBO single-level reformulations have been proposed wiesemann2013pessimistic ; zeng2020practical ; benchouk2024relaxation , their intricate structures continue to pose difficulties for the development of fully first-order gradient-based solution methods. Recently, guanadaprox proposed AdaProx, a gradient-based method for PBO. This AdaProx method employs a double-loop procedure and requires second-order derivative information. This motivates our central research question:
@@
As indicated by the formulation in (1), the PBO can be solved by minimizing the function $\phi(x)$. However, $\phi(x)$ is the value function of a maximization problem whose feasible region depends on the solution set of another optimization problem. Consequently, $\phi(x)$ is generally non-smooth guo2024sensitivity , and evaluating its value and gradient (or subgradient) poses significant computational challenges. The non-smoothness of $\phi(x)$ constitutes a primary challenge in solving PBO, rendering the direct minimization of $\phi(x)$ difficult.
@@ 1.2 Related work
Optimistic Bilevel Optimization: OBO has been extensively studied, with surveys detailing its theory, algorithms, and applications colson2007overview ; dempe2013bilevel ; dempe2020bilevel . A common approach for solving OBO is to reduce it to a single-level problem, using Karush-Kuhn-Tucker (KKT) conditions, leading to Mathematical Programs with Complementarity Constraints (MPCC) allende2013solving ; luo1996mathematical , or through value-function-based inequality constraints ye1995optimality ; outrata1990numerical . Approximating the lower-level solution with a finite trajectory is another strategy maclaurin2015gradient ; pmlr-v70-franceschi17a . These approaches have yielded scalable and efficient algorithms suitable for large-scale machine learning tasks pedregosa2016hyperparameter ; franceschi2018bilevel ; lorraine2020optimizing ; liu2020generic ; liu2021towards ; sow2022convergence ; hong2023two ; arbel2021amortized ; lu2023slm ; ji2021bilevel ; shen2023penalty ; kwon2023fully ; liu2022bome ; lu2024first ; kwon2023penalty . However, the max-structure at the upper level of PBO creates a more complex, three-level-like structure (min-max-min), hindering the direct application of OBO algorithms to PBO.
@@
Pessimistic Bilevel Optimization: PBO has been surveyed in liu2018pessimistic ; dempe2020bilevel . Theoretical studies include aboussoror2001existence , which investigates sufficient conditions for the existence of optimal solutions, and loridan1988approximate , which studies properties of approximate solutions. Optimality conditions for PBO have been explored, including KKT-type conditions for smooth and non-smooth cases dempe2014necessary ; dempe2019two . aussel2019pessimistic studies the relationship between PBO and its MPCC reformulation. For PBO algorithms, aboussoror2005weak ; zheng2013exact propose penalty methods for solving weak linear PBO problems. wiesemann2013pessimistic introduces a semi-infinite programming reformulation of PBO. lampariello2019standard reformulates PBO as an OBO problem with a two-follower Nash game, solving it as an MPCC. zeng2020practical transforms PBO into a minimax problem with coupled constraints, proposing methods for the linear case. More recently, benchouk2024relaxation ; benchouk2025scholtes explores relaxation methods for solving PBO’s KKT conditions. guanadaprox combines the lower-level value function with the KKT conditions of the upper-level max problem, resulting in a constrained minimization problem solved by a gradient-based method. Several heuristic algorithms have also been proposed, though without convergence guarantees alves2018semivectorial ; alekseeva2017matheuristic . Recently, gradient-based algorithms for minimax bilevel optimization have been developed gu2021nonconvex ; hustochastic ; yang2024first . However, these problems differ from PBO in that their max structure is on the upper-level variable, not the lower-level variable, making these algorithms inapplicable to PBO. To our knowledge, fully first-order, single-loop gradient-based algorithms for solving PBO remain limited.
@@ 2.1 Smooth approximation of $\phi(x)$
This reformulation was explored in zeng2020practical as an application of the value function approach for designing numerical methods for PBO problem. The equality in (2) is justified by (zeng2020practical, , Lemmas 1, 2); for completeness, a proof is provided in Appendix B.1.
@@
It is important to highlight the role of the coupling term $\langle y,z\rangle$ in (3), introduced alongside the regularization term $\frac{\sigma}{2}\|z\|^{2}$. This coupling term is crucial for establishing Lemma 2.2, which is the foundation of the asymptotic convergence of the proposed approximation $\phi_{\rho,\sigma}(x)$ to $\phi(x)$ and the sequential convergence of $(y^{*}(x_{k}),z^{*}(x_{k}))$ established in Theorem 2.5 and2.6.
@@ 2.2 Asymptotic convergence of the approximation
Lower semi-continuity is equivalent to the closedness of the function’s epigraph and its level sets, and it guarantees the existence of a minimizer for $\phi(x)$ over a compact set $X$ (see, e.g., [rockafellar2009variational , Theorem 1.9]). Sufficient conditions for Assumption 3, such as the inner semi-continuity of the lower-level solution map $\mathcal{S}(x)$, are discussed in Appendix B.5. Under this assumption, we can establish the following result for epi-convergence.
@@
Conditions (7) (applied with a constant sequence $x_{k}=x$) and (9) together imply the epi-convergence of the sequence of functions $\{\phi_{\rho_{k},\sigma_{k}}\}$ to $\phi$ on $X$ as $k\rightarrow\infty$ (see, e.g., (rockafellar2009variational, , Proposition 7.2)). This signifies that the epigraph of $\phi_{\rho_{k},\sigma_{k}}(x)$ converges, in the set-theoretic sense, to the epigraph of $\phi(x)$. Leveraging this epi-convergence property, and employing results such as (bonnans2013perturbation, , Proposition 4.6) or (rockafellar2009variational, , Theorem 7.31), we can establish the subsequential convergence of minimizers of problem (6).
@@
Beyond the convergence of $\{x_{k}\}$, we establish the convergence behavior of the saddle point $y_{\rho_{k},\sigma_{k}}^{*}(x_{k})$ and $z_{\rho_{k},\sigma_{k}}^{*}(x_{k})$ below, demonstrating that our method yields solutions to both upper and lower level variables.
@@
Assume $\phi(x)$ is lower semi-continuous on $X$. Let $\{\rho_{k}\}$ and $\{\sigma_{k}\}$ be sequences such that $\rho_{k}\rightarrow\infty$ and $\sigma_{k}\rightarrow 0$ as $k\rightarrow\infty$. For any sequence $\{x_{k}\}\subset X$ converging to $\bar{x}$, we have:
@@ 3 Single-loop gradient-based algorithm
Based on the preceding components, we now formally present the Single-loop Pessimistic Bilevel Algorithm (SiPBA) for solving the PBO problem (1) in Algorithm 1. With the assumption that the projections on $X$ and $Y$ are computationally efficient, SiPBA is a single loop fully first order algorithm.
@@ 4 Convergence analysis
$\lim_{k\rightarrow\infty}\|x^{k}-\mathrm{Proj}_{X}\left(x^{k}-\alpha_{k}\nabla\phi_{k}(x^{k})\right)\|/\alpha_{k}=0,\quad\text{and}\quad\lim_{k\rightarrow\infty}\|u^{k}-u_{k}^{*}(x^{k})\|=0.$
The selection of the parameters in Theorem 4.2 is discussed in detail in Appendix A.4. Additionally, we can establish a modified stationarity result for the iterates $x^{k}$ generated by SiPBA in terms of the $\epsilon$-subdifferential (cf. (mordukhovich2018variational, , Theorem 1.26)).
@@
Assume $\phi(x)$ is lower semi-continuous on $X$. Let $\{(x^{k},y^{k},z^{k})\}$ be the sequence generated by SiPBA(Algorithm 1) with parameters chosen as specified in Theorem 4.2. Suppose that the function $\phi(x)$ is bounded below on the set $X$. If $\beta_{0}/\sigma_{0}$ is sufficiently small, then for any $\epsilon>0$ and $\tilde{\epsilon}>0$, there exists an integer $K_{0}>0$ such that for all $k\geq K_{0}$, there exists $\delta_{k}>0$ for which the following holds:
$\phi(x)+\epsilon\|x-x^{k}\|\geq\phi(x^{k})-\tilde{\epsilon},\qquad\forall x\in\mathbb{B}_{\delta_{k}}(x^{k})\cap X.$
@@ 5.1 Synthetic example
We compare SiPBA against two other gradient-based methods—AdaProx-PD and AdaProx-SG guanadaprox —as well as two MPCC-based approaches— Compact Scholtes (Scholtes-C) and Detailed Scholtes (Scholtes-D) relaxation method benchouk2025scholtes . SiPBA, AdaProx-PD, and AdaProx-SG are run for 20,000 iterations, while Scholtes-C and Scholtes-D are run for 10 outer iterations (as they converge within this range). We report the minimum and maximum relative errors, the number of successful runs achieving the tolerance $\epsilon_{rel}<10^{-4}$ (Valid Runs), and the average runtime to reach this tolerance for those valid runs (Ave. Time). Figure 1(a) shows the convergence curve of the gradient-based algorithms and Table 1 summarizes the final performance metrics of all the methods. We further evaluate SiPBA’s robustness to hyperparameters (stepsizes $\alpha_{0},\beta_{0}$ and update factors $p,q,s$) and its scalability by measuring runtime and iterations required to achieve the tolerance, $\epsilon_{rel}<10^{-4}$, across varying hyperparameters and problem dimensions, with results shown in Table 5.1 and Figure 1 (b). All the results indicate the consistent performance and computational efficiency of SiPBA.
@@ 5.2 Spam classification
Spam classification is challenging due to adversarial dynamics and poor cross-domain generalization. We consider the PBO model for spam classification tasks, as proposed by bruckner2011stackelberg :
@@
We conduct a two-part empirical comparison. First, we compare the PBO model (17) trained using SiPBA (with $\varphi(x)$ as the top $k$ principal components) to the same model trained using the SQP method with $\varphi(x)=x$, as proposed in bruckner2011stackelberg . Second, we compare the SiPBA-trained PBO model against a standard single-level model, $\min_{w}l(w,x,y)+\lambda_{1}\text{Reg}(w)$, trained using the scikit-learn library pedregosa2011scikit . We use either hinge loss or cross-entropy for both $l$ and $l^{\prime}$, and refer to the resulting methods as SiPBA-Hinge/CE, SQP-Hinge/CE, and Single-Hinge/CE.
Experiments are conducted using four standard spam datasets: TREC2006 ounis2006overview , TREC2007 macdonald2007overview , EnronSpam metsis2006spam , and LingSpam androutsopoulos2000evaluation . The average results over ten independent runs are summarized in Table 2, which indicate that the PBO model (either trained with SQP or SiPBA) exhibits superior cross-domain performance compared to the single-level models. Moreover, the SiPBA-trained models achieve the best overall accuracy and F1 score.

### v4

@@ 1 Introduction
where $x\in\mathbb{R}^{n}$ and $y\in\mathbb{R}^{m}$ represent the upper-level and lower-level decision variables, respectively, and ${X}\subseteq\mathbb{R}^{n}$ and ${Y}\subseteq\mathbb{R}^{m}$ are closed convex sets. The functions $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ and $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ are the upper-level and lower-level objective functions, respectively. Bilevel optimization naturally models non-cooperative game between two players, often referred to as a Stackelberg game von. Stackelberg (1952). When the lower‐level problem in bilevel optimization admits multiple optimal solutions for a given $x$, the corresponding decision variable $y$ in the upper‐level objective becomes ambiguous. To resolve this, bilevel optimization is commonly formulated in two distinct settings: Optimistic Bilevel Optimization (OBO) and Pessimistic Bilevel Optimization (PBO).
@@
In recent years, bilevel optimization has garnered substantial interest within the machine learning community, finding applications in areas such as hyperparameter optimization Franceschi et al. (2017a), adversarial learning Zhang et al. (2022), reinforcement learning Zheng and Gu (2025), and meta-learning Gao and Sener (2020), among others, where first-order gradient-based methods are preferred for their efficiency and scalabilityBottou et al. (2018); Lin et al. (2020); Sra et al. (2011).
Much of the existing bilevel research focuses on the OBO case, for which numerous fully first-order gradient-based algorithms suitable for large-scale machine learning tasks have been developed Kwon et al. (2023); Liu et al. (2022, 2024); Shen and Chen (2023), often by reformulating the problem using Karush-Kuhn-Tucker (KKT) conditions or through value-function-based constraints. While OBO benefits from a well-established algorithmic toolkit, PBO remains comparatively underexplored from an algorithmic standpoint. PBO offers a robust framework for leaders concerned with worst-case follower responses and a growing body of work has explored the potential of PBO in various machine learning applications such as adversarial learning Brückner and Scheffer (2011); Benfield et al. (2024), contextual optimization Bucarey et al. (2024); Jiménez et al. (2025) and hyperparameter optimization Ustun et al. (2024). Outside the machine learning domain, PBO has found applications in many other practical scenarios, including but not limited to demand response managementKis et al. (2021), rank pricing and second-best toll pricing Calvete et al. (2024); Ban et al. (2009), production-distribution planning Zheng et al. (2016), and gene knockout modelZeng (2020). Yet, the inherent max-structure at the upper level of PBO creates a more complex, three-level-like structure (min-max-min), making the direct application of gradient-based techniques developed for OBO challenging. Although several PBO single-level reformulations have been proposed Wiesemann et al. (2013); Zeng (2020); Benchouk et al. (2024) , their intricate structures continue to pose difficulties for the development of fully first-order gradient-based solution methods. Recently, Guan et al. (2025) proposed AdaProx, a gradient-based method for PBO. This AdaProx method employs a double-loop procedure and requires second-order derivative information. This motivates our central research question:
@@
As indicated by the formulation in (1), the PBO can be solved by minimizing the function $\phi(x)$. However, $\phi(x)$ is the value function of a maximization problem whose feasible region depends on the solution set of another optimization problem. Consequently, $\phi(x)$ is generally non-smooth Guo et al. (2024), and evaluating its value and gradient (or subgradient) poses significant computational challenges. The non-smoothness of $\phi(x)$ constitutes a primary challenge in solving PBO, rendering the direct minimization of $\phi(x)$ difficult.
@@ 1.2 Related work
Optimistic Bilevel Optimization: OBO has been extensively studied, with surveys detailing its theory, algorithms, and applications Colson et al. (2007); Dempe and Zemkoho (2013, 2020). A common approach for solving OBO is to reduce it to a single-level problem, using Karush-Kuhn-Tucker (KKT) conditions, leading to Mathematical Programs with Complementarity Constraints (MPCC) Allende and Still (2013); Luo et al. (1996), or through value-function-based inequality constraints Ye and Zhu (1995); Outrata (1990). Approximating the lower-level solution with a finite trajectory is another strategy Maclaurin et al. (2015); Franceschi et al. (2017b). These approaches have yielded scalable and efficient algorithms suitable for large-scale machine learning tasks Pedregosa (2016); Franceschi et al. (2018); Lorraine et al. (2020); Liu et al. (2020, 2021); Sow et al. (2022); Hong et al. (2023); Arbel and Mairal (2022); Lu (2023); Ji et al. (2021); Shen and Chen (2023); Kwon et al. (2023); Liu et al. (2022); Lu and Mei (2024); Kwon et al. (2024). However, the max-structure at the upper level of PBO creates a more complex, three-level-like structure (min-max-min), hindering the direct application of OBO algorithms to PBO.
@@
Pessimistic Bilevel Optimization: PBO has been surveyed in Liu et al. (2018); Dempe and Zemkoho (2020). Theoretical studies include Aboussoror and Loridan (2001), which investigates sufficient conditions for the existence of optimal solutions, and Loridan and Morgan (1988), which studies properties of approximate solutions. Optimality conditions for PBO have been explored, including KKT-type conditions for smooth and non-smooth cases Dempe et al. (2014, 2019). Aussel and Svensson (2019) studies the relationship between PBO and its MPCC reformulation. For PBO algorithms, Aboussoror and Mansouri (2005); Zheng et al. (2013) propose penalty methods for solving weak linear PBO problems. Wiesemann et al. (2013) introduces a semi-infinite programming reformulation of PBO. Lampariello et al. (2019) reformulates PBO as an OBO problem with a two-follower Nash game, solving it as an MPCC. Zeng (2020) transforms PBO into a minimax problem with coupled constraints, proposing methods for the linear case. More recently, Benchouk et al. (2024, 2025) explores relaxation methods for solving PBO’s KKT conditions. Guan et al. (2025) combines the lower-level value function with the KKT conditions of the upper-level max problem, resulting in a constrained minimization problem solved by a gradient-based method. Several heuristic algorithms have also been proposed, though without convergence guarantees Alves and Antunes (2018); Alekseeva et al. (2017). Recently, gradient-based algorithms for minimax bilevel optimization have been developed Gu et al. (2021); Hu et al. (2021); Yang et al. (2024). However, these problems differ from PBO in that their max structure is on the upper-level variable, not the lower-level variable, making these algorithms inapplicable to PBO. To our knowledge, fully first-order, single-loop gradient-based algorithms for solving PBO remain limited.
@@ 2.1 Smooth approximation of $\phi(x)$
This reformulation was explored in Zeng (2020) as an application of the value function approach for designing numerical methods for PBO problem. The equality in (2) is justified by (Zeng, 2020, Lemmas 1, 2); for completeness, a proof is provided in Appendix B.1.
@@
It is important to highlight the role of the coupling term $\langle y,z\rangle$ in (3), introduced alongside the regularization term $\frac{\sigma}{2}\|z\|^{2}$. This coupling term is crucial for establishing Lemma 2.2, which is the foundation of the asymptotic convergence of the proposed approximation $\phi_{\rho,\sigma}(x)$ to $\phi(x)$, and of the saddle point $(y_{\rho,\sigma}^{*}(x),z_{\rho,\sigma}^{*}(x))$ as established in Theorems 2.5 and 2.6, respectively.
@@ 2.2 Asymptotic convergence of the approximation
Lower semi-continuity is equivalent to the closedness of the function’s epigraph and its level sets, and it guarantees the existence of a minimizer for $\phi(x)$ over a compact set $X$ (see, e.g., (Rockafellar and Wets, 2009, Theorem 1.9)). Sufficient conditions for Assumption 3, such as the inner semi-continuity of the lower-level solution map $\mathcal{S}(x)$, are discussed in Appendix B.5. Under this assumption, we can establish the following result for epi-convergence.
@@
Conditions (7) (applied with a constant sequence $x_{k}=x$) and (9) together imply the epi-convergence of the sequence of functions $\{\phi_{\rho_{k},\sigma_{k}}\}$ to $\phi$ on $X$ as $k\rightarrow\infty$ (see, e.g., (Rockafellar and Wets, 2009, Proposition 7.2)). This signifies that the epigraph of $\phi_{\rho_{k},\sigma_{k}}(x)$ converges, in the set-theoretic sense, to the epigraph of $\phi(x)$. Leveraging this epi-convergence property, and employing results such as (Bonnans and Shapiro, 2013, Proposition 4.6) or (Rockafellar and Wets, 2009, Theorem 7.31), we can establish the subsequential convergence of minimizers of problem (6).
@@
In the following, we characterize the asymptotic behavior of the saddle point $(y_{\rho,\sigma}^{*}(x),z_{\rho,\sigma}^{*}(x))$ as $\rho\rightarrow\infty$ and $\sigma\rightarrow 0$, and show that both components converge to the solution of the maximization problem that defines the value function $\phi(x)$ in (1).
@@
Assume $\phi(x)$ is lower semi-continuous on $X$. Let $\{\rho_{k}\}$ and $\{\sigma_{k}\}$ be sequences such that $\rho_{k}\rightarrow\infty$ and $\sigma_{k}\rightarrow 0$ as $k\rightarrow\infty$ and let $\{x_{k}\}$ be a sequence such that $x_{k}\in X$ and $x_{k}\rightarrow\bar{x}$ as $k\rightarrow\infty$. Then, we have:
@@ 3 Single-loop gradient-based algorithm
Based on the preceding components, we now formally present the Single-loop Pessimistic Bilevel Algorithm (SiPBA) for solving the PBO problem (1) in Algorithm 1. In many practical applications where projections onto $X$ and $Y$ are computationally efficient, SiPBA offers the significant advantage of a single-loop structure, making it straightforward to implement.
@@ 4 Convergence analysis
$\underset{k\rightarrow\infty}{\lim\inf}\;\|x^{k}-\mathrm{Proj}_{X}\left(x^{k}-\alpha_{k}\nabla\phi_{k}(x^{k})\right)\|/\alpha_{k}=0,\quad\text{and}\quad\underset{k\rightarrow\infty}{\lim\inf}\;\|u^{k}-u_{k}^{*}(x^{k})\|=0.$
Based on Theorem 4.2, a practical parameter selection strategy for SiPBA is provided in Appendix A.4. Furthermore, we can establish a modified stationarity result for the iterates $x^{k}$ generated by SiPBA in terms of the $\epsilon$-subdifferential (cf. (Mordukhovich, 2018, Theorem 1.26)).
@@
Assume $\phi(x)$ is lower semi-continuous on $X$. Let $\{(x^{k},y^{k},z^{k})\}$ be the sequence generated by SiPBA(Algorithm 1) with parameters chosen as specified in Theorem 4.2. Suppose that the function $\phi(x)$ is bounded below on the set $X$. If $\beta_{0}/\sigma_{0}$ is sufficiently small, then there exists a subsequence $\{x^{k_{j}}\}$ such that for any $\epsilon>0$ and $\tilde{\epsilon}>0$, there exists an integer $K>0$ such that for all $k_{j}>K$, there exists a corresponding $\delta_{j}>0$ for which the following inequality holds:
$\phi(x)+\epsilon\|x-x^{k_{j}}\|\geq\phi(x^{k_{j}})-\tilde{\epsilon},\qquad\forall x\in\mathbb{B}_{\delta_{j}}(x^{k_{j}})\cap X.$
@@ 5.1 Synthetic example
We compare SiPBA against two other gradient-based methods—AdaProx-PD and AdaProx-SG Guan et al. (2025)—as well as two MPCC-based approaches— Compact Scholtes (Scholtes-C) and Detailed Scholtes (Scholtes-D) relaxation method Benchouk et al. (2025). SiPBA, AdaProx-PD, and AdaProx-SG are run for 20,000 iterations, while Scholtes-C and Scholtes-D are run for 10 outer iterations (as they converge within this range). We report the minimum and maximum relative errors, the number of successful runs achieving the tolerance $\epsilon_{rel}<10^{-4}$ (Valid Runs), and the average runtime to reach this tolerance for those valid runs (Ave. Time). Figure 1(a) shows the convergence curve of the gradient-based algorithms and Table 1 summarizes the final performance metrics of all the methods. We further evaluate SiPBA’s robustness to hyperparameters (stepsizes $\alpha_{0},\beta_{0}$ and update factors $p,q,s$) and its scalability by measuring runtime and iterations required to achieve the tolerance, $\epsilon_{rel}<10^{-4}$, across varying hyperparameters and problem dimensions, with results shown in Table 5.1 and Figure 1 (b). All the results indicate the consistent performance and computational efficiency of SiPBA.
@@ 5.2 Spam classification
Spam classification is challenging due to adversarial dynamics and poor cross-domain generalization. We consider the PBO model for spam classification tasks, as proposed by Brückner and Scheffer (2011):
@@
We conduct a two-part empirical comparison. First, we compare the PBO model (17) trained using SiPBA (with $\varphi(x)$ as the top $k$ principal components) to the same model trained using the SQP method with $\varphi(x)=x$, as proposed in Brückner and Scheffer (2011). Second, we compare the SiPBA-trained PBO model against a standard single-level model, $\min_{w}l(w,x,y)+\lambda_{1}\text{Reg}(w)$, trained using the scikit-learn library Pedregosa et al. (2011). We use either hinge loss or cross-entropy for both $l$ and $l^{\prime}$, and refer to the resulting methods as SiPBA-Hinge/CE, SQP-Hinge/CE, and Single-Hinge/CE.
Experiments are conducted using four standard spam datasets: TREC2006 Ounis et al. (2006), TREC2007 Macdonald et al. (2007), EnronSpam Metsis et al. (2006), and LingSpam Androutsopoulos et al. (2000). The average results over ten independent runs are summarized in Table 2, which indicate that the PBO model (either trained with SQP or SiPBA) exhibits superior cross-domain performance compared to the single-level models. Moreover, the SiPBA-trained models achieve the best overall accuracy and F1 score.